What are the values for column a and column b. Assume successive crop rotation and discount rate is 3%. Please show at least one of the calculations for each column.
Stand age |Timber |value volume per m Annual |a) Mean |Site valu Increment 20 65 70 75 80 85 90 1535 1755 1950 2111 2235 2340 24 26 28 30
Answer #1

Mean Annual Increment = Total Volume / Age

Mean Annual Increment at Standing Age 65 = 1535 /65 = 23.62 meter cube

Mean Annual Increment at Standing Age 70 = 1755 /70 = 25.07 meter cube

Mean Annual Increment at Standing Age 75 = 1950 /75 = 26 meter cube

Mean Annual Increment at Standing Age 80 = 2111 /80 = 26.39 meter cube

Mean Annual Increment at Standing Age 85 = 2235 /85 = 26.29 meter cube

Mean Annual Increment at Standing Age 90 = 2340 /90 = 26 meter cube

PV of site value = Periodic income / ((1+discount rate)n -1)

PV rotation at the age of 65 = 1535*20 / ((1.03)65 -1) = 30700/(6.83-1) = 30700 / 5.83 = 5266

(1.03)65 =6.829982735

PV rotation at the age of 70 = 1755*22 / ((1.03)70-1) = 38610/(7.91-1) = 38610 / 6.91 = 5588

(1.03)70 = 7.917821912

PV rotation at the age of 75 = 1950*24 / ((1.03)75-1) = 46800/(9.18-1) = 46800 / 8.18= 5721

(1.03)75 = 9.178925668

PV rotation at the age of 80 = 2111*26 / ((1.03)80-1) = 54886 /(10.64-1) = 54886 /9.64= 5694

(1.03)80= 10.64089056

PV rotation at the age of 85 = 2235*28 / ((1.03)85-1) = 62580 /(12.34-1) = 62580 /11.34= 5518

(1.03)85=12.33570855

PV rotation at the age of 90 = 2340*30 / ((1.03)90-1) = 70200 /(14.30-1) = 70200 /13.30= 5278

(1.03)90= 14.30046711